# Cross-modal Ship Re-Identification via Optical and SAR Imagery: A Novel Dataset and Method

### v1

@@ 1 Introduction
Figure 2: Schematic diagram of ship tracking method based on multimodal LEO constellations. By performing ReID on the ship of interest through remote sensing ship detection at different times and locations, the spatiotemporal trajectory of the ship can be established. This allows for the prediction of the ship’s future travel direction, enabling the constellation to be guided for the next imaging.
@@ 2 Related Work
Video satellites have also been employed for ship tracking. Li et al. [19] released single-object tracking datasets for ships using satellite video, while Yin et al. [50] and Li et al. [17] introduced multi-object tracking datasets. These video satellites typically offer a ground sample distance (GSD) of about 1 meter, with frame rates ranging from 5 to 30 frames per second, and capture footage lasting between 60 to 120 seconds [16, 19]. Researchers have developed specialized methods tailored to these datasets, achieving precise single-object or multi-object ship tracking [44, 18, 52, 56]. The primary technical challenge for these approaches lies in ship detection, as the tracking itself is relatively straightforward. However, due to the motion of video satellites, tracking is often restricted to short durations of just a few tens of seconds. Although these methods can provide valuable information on the ship’s course, speed, and other dynamic attributes, achieving long-term tracking necessitates the integration of other satellite constellations, along with robust identity recognition and data association techniques. Lang et al. [11] made initial attempts at ship tracking using SAR constellations, employing a detection-matching-tracking strategy that demonstrated the method’s feasibility. However, their work was limited by insufficient data and lacked further experimental validation.
@@
Cross-modal ReID. Cross-modal ReID is commonly employed when RGB images alone are insufficient, and additional modalities, such as infrared or sketches, can provide complementary information. Among these, infrared-visible pedestrian ReID has seen the most significant progress. Several cross-modal pedestrian ReID datasets have been introduced, including SYSU-MM01 [39] and RegDB [49], which have become widely accepted as standard benchmarks in the field. Li et al. [12] proposed a method to bridge the gap between RGB and infrared images by embedding RGB images into an X-modality space using a lightweight network architecture. Wu et al. proposed MPANet [40], incorporating a modality alleviation module and a pattern alignment module to discover cross-modal nuances in different patterns. In another study, Park et al. [29] addressed cross-modal discrepancies by leveraging dense correspondences between images of individuals across different modalities. Further advancing this area, Wu et al. [41] proposed an unsupervised method based on progressive graph matching and cross contrastive learning to extract modality-invariant features, while Zhang et al. [53] introduced an embedding space enhancement network designed to generate more diverse embeddings, thus providing more informative feature representations. In a novel approach, Li et al. [15] utilized a pre-trained large model as an encoder, facilitating effective multimodal ReID across RGB, infrared, sketch, and textual modalities without the need for additional fine-tuning. While infrared and visible light images share more common properties, the differences between optical and SAR images are much greater. These methods struggle to suppress the redundant information between the two modalities, making them less suited for optical-SAR ReID tasks.
@@ 3.1 Dataset Collection
Given the stringent requirements for creating this dataset, publicly accessible data is nearly nonexistent, we opted to acquire the imagery from scratch through programmed satellite imaging. However, this approach still presents several challenges: (1) The high cost of satellite tasking significantly limits imaging opportunities; (2) The dynamic nature of ships at sea makes it difficult to capture the same ship multiple times to ensure consistent tracking; (3) Imaging conditions are constrained by factors such as weather, lighting, and satellite transit times, requiring precisely timed captures to obtain a multimodal image sequence of a single object; (4) Different modalities must be acquired through various constellations, demanding careful coordination of capture windows. As a result, completing the image acquisition for this dataset is a large-scale endeavor that requires substantial time and financial resources.
To mitigate these challenges, we implement several strategies: (1) We prioritize ports, canals, and similar locations as primary imaging sites, enabling each capture to include a high density of ships; (2) We minimize the interval between imaging sessions and concentrate on capturing ships that are anchored, thereby reducing the impact of vessel movement. These approaches enable us to acquire images of the same ship from multiple angles and in different modalities using various satellites, where the ship’s motion or precise location do not adversely affect the imaging results, and makes no difference for the ship ReID task. Some examples of the images we captured are shown in Figure 3.
@@ 3.2 Dataset Description
Table 1: Detailed statistics for HOSS ReID dataset. “(+163)” indicates there are 163 distractor objects added to the gallery.
@@
Our dataset format is modeled after the Market-1501 [59] dataset, with filenames containing the object ID, sequence ID, camera ID, and modality. The sequence ID specifies which of the 13 sequences the image belongs to, while the camera ID indicates which satellite captured the image within that sequence. All image files are in “TIF” format, with optical images formatted as 3-channel 8-bit RGB and SAR images as single-channel 32-bit floating-point to preserve more detailed information.
We divide the images into training and validation sets, with specific image and object counts detailed in Table 1. To enhance the dataset’s complexity and better simulate real-world scenarios, we include 163 distractor objects in the gallery that do not belong to the query, labeled as “-1”. These distractor images share the same origin as other images but correspond to no specific queries, serving to expand the gallery scale and simulate real-world scenarios containing abundant irrelevant objects. Our dataset features queries consisting of both optical and SAR images, each accounting for 50%. However, the number of each modality per object in the gallery is arbitrary. Similarly, in the training set, the distribution of modalities per object is also arbitrary, with a small proportion of objects represented by images of only one modality.
@@ 5 Experiments
Table 2: Comparisons of CMC (%) and mAP (%) performances on the HOSS ReID dataset. The star * indicates TransOSS without pre-training on optical-SAR image pairs. “Optical to SAR” means that queries from the optical modality are searched only within the SAR modality gallery, and “SAR to Optical” has a similar meaning.
@@ 5.3 Cross-modal experiments
To further validate the effectiveness of our method for cross-modal ReID, we employed Grad-CAM[34] to visualize the attention maps of TransOSS, as shown in Figure 6. The visualization highlights the significant morphological differences between ships in the two modalities, but TransOSS leverages a modality-shared transformer encoder to map images from both modalities into a shared feature space, extracting common cross-modal features. The attention maps reveal that the model focuses on similar local areas and ship contours across both modalities, demonstrating that it has learned to capture both modality-invariant details and global features, such as the ship’s shape.
@@ 5.4 Ablation Study
Table 3: The ablation study of TransOSS. “CDT” represents cross-modal dual-head tokenizer, “MIE” represents modality information embeddings, and “SSE” represents ship size embeddings in the table.
@@ 5.5 Discussions
Unlike conventional images, remote sensing imagery offers physically meaningful information, such as geographic coordinates and the physical size of objects, and SAR images further reflect physical characteristics, such as specific structures that generate scattering points, providing valuable additional information that can be effectively utilized. The remote sensing field already offers a range of rich datasets, some of which could supplement network training for this task and potentially yield strong results. Alternatively, self-supervised or unsupervised methods [5, 26, 1] should be explored, alongside the incorporation of additional data augmentation techniques [6, 10], which are particularly effective in data-scarce scenarios. Future work could also expand this task by incorporating additional modalities, such as textual information and multispectral imagery, to broaden its applicability across a wider range of scenarios.

### v2

@@ 1 Introduction
Figure 2: Schematic diagram of ship tracking method based on multimodal LEO constellations. By performing ReID on the ship of interest through remote sensing ship detection at different times and locations, the spatiotemporal trajectory of the ship can be established. This enables predicting the ship’s trajectory to guide subsequent satellite imaging.
@@ 2 Related Work
Video satellites have also been employed for ship tracking. Li et al. [19] released single-object tracking datasets for ships using satellite video, while Yin et al. [50] and Li et al. [17] introduced multi-object tracking datasets. These video satellites typically offer a ground sample distance (GSD) of about 1 meter, with frame rates ranging from 5 to 30 frames per second, and capture footage lasting between 60 to 120 seconds [16, 19]. Researchers have developed specialized methods tailored to these datasets, achieving precise single-object or multi-object ship tracking [44, 18, 52, 56]. The main technical challenge lies in ship detection, as tracking is relatively straightforward but limited to short durations (tens of seconds) due to satellite motion. While these methods reveal valuable dynamic attributes like course and speed, long-term tracking requires integrating other constellations and employing robust identity recognition and data association techniques. Lang et al. [11] made initial attempts at ship tracking using SAR constellations, employing a detection-matching-tracking strategy that demonstrated the method’s feasibility. However, their work was limited by insufficient data and lacked further experimental validation.
@@
Cross-modal ReID. Cross-modal ReID is commonly employed when RGB images alone are insufficient, and additional modalities, such as infrared or sketches, can provide complementary information. Among these, infrared-visible pedestrian ReID has seen the most significant progress. Several cross-modal pedestrian ReID datasets have been introduced, including SYSU-MM01 [39] and RegDB [49], which have become widely accepted as standard benchmarks in the field. Li et al. [12] proposed a method to bridge the gap between RGB and infrared images by embedding RGB images into an X-modality space using a lightweight network architecture. Wu et al. proposed MPANet [40], incorporating a modality alleviation module and a pattern alignment module to discover cross-modal nuances in different patterns. In another study, Park et al. [29] addressed cross-modal discrepancies by leveraging dense correspondences between images of individuals across different modalities. Further advancing this area, Wu et al. [41] proposed an unsupervised method based on progressive graph matching and cross contrastive learning to extract modality-invariant features, while Zhang et al. [53] introduced an embedding space enhancement network designed to generate more diverse embeddings, thus providing more informative feature representations. In a novel approach, Li et al. [15] utilized a pre-trained large model as an encoder, facilitating effective multimodal ReID across RGB, infrared, sketch, and textual modalities without the need for additional fine-tuning. Compared to infrared and visible images, optical and SAR images differ significantly, making it difficult for existing methods to suppress cross-modal redundancy, thus limiting their effectiveness in optical-SAR ReID tasks.
@@ 3.1 Dataset Collection
Given the stringent requirements for creating this dataset, publicly accessible data is nearly nonexistent, we opted to acquire the imagery from scratch through programmed satellite imaging. However, this approach faces several challenges: (1) the high cost of satellite tasking limits imaging opportunities; (2) the mobility of ships hinders repeated observations; (3) imaging conditions are constrained by weather, lighting, and satellite transit times, necessitating precise timing to capture multimodal sequences.
To mitigate these challenges, we implement several strategies: (1) We prioritize ports, canals, and similar locations as primary imaging sites, enabling each capture to include a high density of ships; (2) We minimize the interval between imaging sessions and concentrate on capturing ships that are anchored, thereby reducing the impact of vessel movement, where the ship’s motion or precise location do not adversely affect the imaging results, and makes no difference for the ship ReID task.. Some examples of the images we captured are shown in Figure 3.
@@ 3.2 Dataset Description
Figure 4: (a) Distribution of ship slice heights. (b) Dataset construction process. (c) Distribution of ship categories.
@@
Table 1: Detailed statistics for HOSS ReID dataset. “(+163)" indicates there are 163 distractor objects added to the gallery.
@@
Our dataset format is modeled after the Market-1501 [59] dataset, with filenames containing the object ID, sequence ID, camera ID, and modality. The sequence ID specifies which of the 13 sequences the image belongs to, while the camera ID indicates which satellite captured the image within that sequence. All image files are in “TIF" format, with optical images formatted as 3-channel 8-bit RGB and SAR images as single-channel 32-bit floating-point to preserve more detailed information. Some statistical information of the dataset is presented in Figure 4: the length distribution of images (reflecting the size distribution of ships), the dataset construction process, and the general category distribution.
We divide the images into training and validation sets, with specific image and object counts detailed in Table 1. To enhance the dataset’s complexity and better simulate real-world scenarios, we include 163 distractor objects in the gallery that do not belong to the query, labeled as “-1". These distractor images share the same origin as other images but correspond to no specific queries, serving to expand the gallery scale and simulate real-world scenarios containing abundant irrelevant objects. Our dataset features queries consisting of both optical and SAR images, each accounting for 50%. However, the number of each modality per object in the gallery is arbitrary. Similarly, in the training set, the distribution of modalities per object is also arbitrary, with a small proportion of objects represented by images of only one modality.
@@ 5 Experiments
Table 2: Comparisons of CMC (%) and mAP (%) performances on the HOSS ReID dataset. The star * indicates TransOSS without pre-training on optical-SAR image pairs. “Optical to SAR" means that queries from the optical modality are searched only within the SAR modality gallery, and “SAR to Optical" has a similar meaning.
@@ 5.3 Cross-modal experiments
To further validate the effectiveness of our method for cross-modal ReID, we employed Grad-CAM [34] to visualize the attention maps of TransOSS, as shown in Figure 7. They reveal that the model focuses on similar local areas and ship contours across both modalities, demonstrating that it has learned to capture both modality-invariant details and global features, such as the ship’s shape.
@@ 5.4 Ablation Study
Table 3: The ablation study of TransOSS. “CDT" represents cross-modal dual-head tokenizer, “MIE" represents modality information embeddings, and “SSE" represents ship size embeddings in the table.
@@ 5.5 Discussions
Unlike conventional images, remote sensing imagery offers physically meaningful information, such as geographic coordinates and the physical size of objects, and SAR images further reflect physical characteristics, such as specific structures that generate scattering points, providing valuable additional information that can be utilized. Alternatively, self-supervised or unsupervised methods [5, 26, 1] should be explored, alongside the incorporation of additional data augmentation techniques [6, 10], which are particularly effective in data-scarce scenarios. Future work could also incorporate modalities like text and multispectral imagery to enhance applicability across diverse scenarios.
@@ 6 Conclusion
Acknowledgement
This work was supported by the Key Deployment Program of the Chinese Academy of Sciences (Grant No. KGFZD-145-23-18).